OpenStudy (anonymous):

v=(pie)r^2h solve for h

OpenStudy (anonymous):

h = v/(Pi*r^2)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

ı dont think it is true

OpenStudy (anonymous):

just divide both sides by (pie)r^2

OpenStudy (anonymous):

v/pi=r^2h lnv/pi=2hlnr ln(v/pi-r)/2=h

OpenStudy (anonymous):

no v=(pie)r^2h v/((pie)r^2)=h divide both sides by (pie)r^2 It is not r^(2h) It is r^2 * h

OpenStudy (anonymous):

I think it is volume of a cylinder

OpenStudy (anonymous):

all i had to do was solve for h and yes it is the volume for a cylinder.